OpenStudy (anonymous):

I got this question wrong on my homework.. Can anyone show me step by step how to do this please? (-2x^2/5)^2

OpenStudy (unklerhaukus):

\[\left(\frac{-2x^2}{5}\right)^2\] \[=\frac{\left(-2x^2\right)^2}{5^2}\]

OpenStudy (anonymous):

\[\left( \frac{-2x^2}{5}\right)^2=\frac{-2x^2}{5} \cdot \frac{-2x^2}{5}\] and then..?

OpenStudy (unklerhaukus):

\[=\left(\frac{(-2\times-2)(x^2\times x^2)}{5\times 5 }\right)\]

OpenStudy (anonymous):

@UnkleRhaukus \[4x ^{4}/ 25\] ?

OpenStudy (unklerhaukus):

right.

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

@UnkleRhaukus So when you have the exponent on the outside is it adding the exponents or multiplying ?

OpenStudy (anonymous):

Multiplying. \[(a^b)^c=a^{bc}\]

OpenStudy (unklerhaukus):

\[(x^n)^m=x^{n\times m}=x^{nm}\]

OpenStudy (anonymous):

WOOOOOOO! Ok thanks boys for all your help! :) Love it! @Limitless @UnkleRhaukus

OpenStudy (anonymous):

You're welcome.

OpenStudy (anonymous):

@Limitless @UnkleRhaukus So if I have \[(4z ^{3}/3)^{2}\] would it be \[16z ^{6}/6\]

OpenStudy (unklerhaukus):

not quite

OpenStudy (anonymous):

I mean over 9 not 6

OpenStudy (anonymous):

Yup.

OpenStudy (unklerhaukus):

that is it

OpenStudy (anonymous):

I forgot to do 3x3 I did 2x3

OpenStudy (anonymous):

lol It confuses me!!!! :(

OpenStudy (anonymous):

It just takes practice, that is all.

OpenStudy (anonymous):

Thanks for your help sweetheart @Limitless

OpenStudy (anonymous):

No problem. Teaching people is fun if they want to learn.